2. ſi 0 Verify Cayley-Hamilton theorem, Find A ^ 4 when A= 2 1 1 1 1 -1 3 -1 1
1
Expert's answer
2020-12-29T16:01:37-0500

"A=\begin{bmatrix}2&1&1\\1&1&-1\\3&-1&1\end{bmatrix}"

Characteristic polynomial of A,

"det(A-\lambda I)=0"

"\Rightarrow det\left(\begin{bmatrix}2&1&1\\1&1&-1\\3&-1&1\end{bmatrix} - \lambda \begin{bmatrix}1&0&0\\0&1&0\\0&0&1\end{bmatrix}\right)=0"

"\Rightarrow det \left(\begin{bmatrix}2-\lambda&1&1\\1&1-\lambda&-1\\3&-1&1-\lambda\end{bmatrix}\right)=0"

"\Rightarrow (2-\lambda)det\left(\begin{bmatrix}1-\lambda&-1\\-1&1-\lambda\end{bmatrix}\right)"

"-det\left(\begin{bmatrix}1&-1\\3&1-\lambda\end{bmatrix}\right)"

"+det\left(\begin{bmatrix}1&1-\lambda\\3&-1\end{bmatrix}\right)=0"


"\Rightarrow \left(2-\lambda\right)\left(\lambda^2-2\lambda\right)-\left(-\lambda+4\right)+ \left(3\lambda-4\right)=0"

"\Rightarrow \lambda^3-4\lambda^2+8=0"

"\therefore" Characteristic polynomial of A :- "\lambda^3-4\lambda^2+8=0"


Now,

"A^2= A\times A =\begin{bmatrix}2&1&1\\1&1&-1\\3&-1&1\end{bmatrix}\times \begin{bmatrix}2&1&1\\1&1&-1\\3&-1&1\end{bmatrix}"


"=\begin{bmatrix}8&2&2\\0&3&-1\\8&1&5\end{bmatrix}"

and,

"A^3=A^2\times A=\begin{bmatrix}8&2&2\\0&3&-1\\8&1&5\end{bmatrix}\times \begin{bmatrix}2&1&1\\1&1&-1\\3&-1&1\end{bmatrix}"


"=\begin{bmatrix}24&8&8\\0&4&-4\\32&4&12\end{bmatrix}"


Now, by Cayley-Hamilton theorem, putting "\lambda=A" ,

"A^3-4A^2+8I=0_{3\times3}"

"\Rightarrow \begin{bmatrix}24&8&8\\0&4&-4\\32&4&12\end{bmatrix}-4\begin{bmatrix}8&2&2\\0&3&-1\\8&1&5\end{bmatrix}"


"+8\begin{bmatrix}1&0&0\\0&1&0\\0&0&1\end{bmatrix}=0_{3\times3}"


"\Rightarrow\begin{bmatrix}0&0&0\\0&0&0\\0&0&0\end{bmatrix}=0_{3\times3}"


"\therefore" Cayley-Hamilton Theorem Verified.


Now, by Cayley-Hamilton theorem,


"A^3-4A^2+8I=0_{3\times3}"

"\Rightarrow A(A^3-4A^2+8I)=A\times0_{3\times3}"

[ Multiplying both side by matrix "A" ]

"\Rightarrow A^4-4A^3+8A=0_{3\times3}"

"\Rightarrow A^4 = 4A^3-8A"

"=4\begin{bmatrix}24&8&8\\0&4&-4\\32&4&12\end{bmatrix}-8\begin{bmatrix}2&1&1\\1&1&-1\\3&-1&1\end{bmatrix}" "=\begin{bmatrix}80&24&24\\-8&8&-8\\104&24&40\end{bmatrix}"



"\therefore A^4=\begin{bmatrix}80&24&24\\-8&8&-8\\104&24&40\end{bmatrix}"